OpenStudy (shamil98):

6. You have a coupon for your favorite clothing store for $20 off any purchase of more than $50. The store is also running a 25%-off sale on its entire inventory. Let x be the original price, f(x) be the price with the $20 coupon applied, and g(x) be the price with the 25% discount applied. Find the expression for f(x) and g(x) f(x) = x - 20; x >= 50 g(x) = (0.75)(x) c. What would the expression (f o g)(x) represent? d. What would the expression (g o f) (x) represent?

OpenStudy (luigi0210):

Sorry Sham, I gtg do something. But I'd say you might wanna re-check it

OpenStudy (shamil98):

Okay.

OpenStudy (rsadhvika):

g(x) = x(0.75) f(x) = x-20

OpenStudy (rsadhvika):

keep them simple, you teacher wants to explain you the concept of composition after this

OpenStudy (rsadhvika):

Now, if u want to know whats the cost, after applying both discount+coupon we could simply do f(g(x))

OpenStudy (rsadhvika):

g(x) > 50 is the constrain btw

OpenStudy (rsadhvika):

f(x) is the price wid discount applied, that makes f(x) = x-20 if x > 50, other wise f(x) = x

OpenStudy (shamil98):

.75 of f(x) = g(x) because it says its 25% off.

OpenStudy (rsadhvika):

similarly g(x) is the price wid discount applied, that makes g(x) = 0.75

OpenStudy (rsadhvika):

dont mess wid f(x), when constructing g(x)

OpenStudy (rsadhvika):

do them independently

OpenStudy (rsadhvika):

* similarly g(x) is the price wid discount applied, that makes g(x) = 0.75x

OpenStudy (shamil98):

Oh, okay, i get it.

OpenStudy (rsadhvika):

good :) overall, \(f(x) = x ; x <= 50\) \(f(x) = x-20 ; x > 50\) \(g(x) = .75x\)

OpenStudy (rsadhvika):

do they look okays ?

OpenStudy (shamil98):

Could you help with the other parts of the question? I've edited it. Yes it makes more sense because x is the original price .

OpenStudy (rsadhvika):

(f o g)(x) = f(g(x))

OpenStudy (rsadhvika):

look carefully, first we're applying g, then giving it to f

OpenStudy (rsadhvika):

so, first we're doing .75x, then we're doing x-20 that means, first we're doing 25% discount, and on top of that, if that amount is > 50, we're subtracting 20

OpenStudy (shamil98):

It wouldn't be .75x - 20; .75x >= 50?..

OpenStudy (rsadhvika):

in simple terms, (f o g)(x) , represents first applying 25% discount, and on top of that, applying the gift coupon

OpenStudy (rsadhvika):

forget about expression, they dont want it here

OpenStudy (shamil98):

Oh, yeah... mb.

OpenStudy (rsadhvika):

:) can u try next part urself,

OpenStudy (shamil98):

Yeah. for (g o f) (x) represents first applying 20 dollar coupon then the 25% discount?

OpenStudy (rsadhvika):

perfect ! cuz (g o f)(x) = g(f(x)) we observe that, first we're giving x to f, after that, we're giving it to g.

OpenStudy (shamil98):

Thank you :)

OpenStudy (rsadhvika):

btw, can we knw which one saves us money ? :) (g o f)(x) or (f o g)(x)

OpenStudy (shamil98):

I think ( f o g) (x) saves more money because it's based off of percentage.

OpenStudy (rsadhvika):

Yup !!! nice work :)

OpenStudy (rsadhvika):

but careful ! you may miss 20 deduction, if that percentage discount makes the amount < 50

OpenStudy (shamil98):

Yeah, that's true. So it would have to be a higher number that when it's reduced by 25% its greater than or equal to 50.

OpenStudy (rsadhvika):

exactly ! you get the point :)